Investment X offers to pay you $5,500 per year for nine years, whereas Investment Y offers to pay you $8,000 per year for five years.


 a. Calculate the present value for Investments X and Y if the discount rate is 5 percent. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

 b. Calculate the present value for Investments X and Y if the discount rate is 15 percent. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

Present value of investment is equal to the present value of future cash inflows

a.Investment X at 5% = 5,500*PVAF(5%, 9 years)

= 5,500*7.10782

= $39,093.01

Investment Y = 8000*PVAF(5%, 5 years)

= 8,000*4.32948

= $34,635.84

b.Investment X at 15% = 5,500*PVAF(15%, 9 years)

= 5,500*4.77158

= $26,243.69

Investment Y = 8000*PVAF(15%, 5 years)

= 8000*3.35216

= $26,817.28